OpenStudy (bananas1234):

Multiply and simplify: b^4 • b • b^2 • b^3

OpenStudy (hsjkirribilli266199):

"b^10" I guess...

OpenStudy (bananas1234):

How did you get that answer?

OpenStudy (kidrah69):

Add up the exponents

OpenStudy (hsjkirribilli266199):

by adding exponents. I am not sure though...

OpenStudy (bananas1234):

well the exponents together would be 9 then we have b that would make it 10. i think

OpenStudy (kidrah69):

when you have the same base and you are multiplying you will add the exponents

OpenStudy (kidrah69):

so in this case 10 is right :)

OpenStudy (thefurball):

think of b as b^1

OpenStudy (hsjkirribilli266199):

the second term is b^1 but we dont write it

OpenStudy (kidrah69):

10 is right

OpenStudy (bananas1234):

ok, so we would not multiply the exponents

OpenStudy (kidrah69):

you have it correct when you have the same base you will ADD

OpenStudy (bananas1234):

ohhh i see so what if we did not have the same base?

OpenStudy (kidrah69):

then that would be something diffrent that comes after this do you know the exponent rules?

OpenStudy (nincompoop):

laws of exponent \(\huge x^a \times x^b = x^{a+b} \)

OpenStudy (bananas1234):

no

OpenStudy (bananas1234):

ohh i see @nincompoop

OpenStudy (nincompoop):

meaning, you can only add the exponents of the same variables

OpenStudy (kidrah69):

Thats what i said

OpenStudy (kidrah69):

http://www.mathsisfun.com/algebra/exponent-laws.html <---look at this it will help u :)

OpenStudy (nincompoop):

you cannot use the same rule for \(\huge m^a \times p^b \) because m and p are not the same variable, unless it was defined somewhere that \(m = p\)

OpenStudy (bananas1234):

Thanks! @kidrah69

OpenStudy (kidrah69):

http://prntscr.com/6n51we <---laws of exponents no problem :)

OpenStudy (nincompoop):

you're welcome

OpenStudy (kidrah69):

thanks nin

OpenStudy (bananas1234):

thanks!! i think i understand it a bit better now!!

OpenStudy (kidrah69):

yep just review the laws and you'll be good :)

OpenStudy (bananas1234):

Also that is an adorable baby in your pic! @kidrah69 Thanks!!

OpenStudy (kidrah69):

LOL thanks ^_^